# Non-linear anomalous Edelstein response at altermagnetic interfaces

### v2

@@ Abstract
Calculation of non-linear Edelstein responses — Having established that, due to the interplay of RSOC with AM symmetry allows anomalous non-linear EEs, we now determine them in an experimentally relevant parameter range. For the effective in-plane field $h_{y}$ resulting from weak ferromagnetism, we choose $h_{y}=40$ meV as benchmark value [17]. The first effect of the in-plane Zeeman coupling is to lift the degeneracy at $\Gamma$, as shown in Fig. 4(a), in contrast with the simple Rashba system, for which the in-plane magnetic field can only move the degeneracy point. Moreover, as mentioned above, it breaks the symmetry of the system both in the shape of the Fermi surface and in the chiral structure of the spin, as one can see in Fig. 4(b). The resulting Edelstein susceptibility tensor $\chi$ is shown in Fig. 4(c). While $\chi_{xx}$ vanishes, $\chi_{xy}$ depends non-monotonically on chemical potential and reaches its maximum after occupying the first band. Compared to the case $h_{y}=0$, $\chi_{xy}$ grows slower to its maximum, but reaches a value similar to the situation with pure RSOC. Moreover, the in-plane magnetic field is able to induce a substantial out-of-plane anomalous transversal response $\chi_{xz}$ due to the large altermagnetic coupling, with the transversal component reaching larger values than the in-plane response, even larger than the canonical linear RSOC situation.
@@
The non-monotonic behaviour of $\chi_{xz}$ as a function of $\mu$ derives from the competition and complex interplay of the different interactions. At low filling, even an infinitesimal $h_{y}$ breaks the $\bf k\to-\bf k$ symmetry in the integrand in Eq. (4), leading to a characteristic threshold behavior in the response. Increasing $\mu$, the out-of-plane response is first slightly suppressed and later saturates to a constant value in the altermagnetic-dominated regime. Crucially, all three ingredients are required to exhibit the out-of-plane response: the altermagnetic interaction enforces an out-of-plane magnetization, the weak ferromagnetism breaks its otherwise symmetric pattern, and the RSOC induces the spin-momentum locking required for this non-linear Edelstein response.
@@
Summary and outlook — Using a continuum description for AMs based on ab-initio electronic band structure calculations, we have shown that the interplay of altermagnetic coupling with an interfacial RSOC gives rise to the appearance of out-of-plane spin polarizations throughout the Brillouin zone via non-linear EEs. While the in-plane linear response can be strongly suppressed with the band filling, the non-linear Edelstein susceptibilities exhibit strong non-monotonic behaviour. The non-linear responses we determined involve a term that depends on the in-plane electric and the magnetic fields, and a second-order term in the electric field. For the former term, an anomalous out-of-plane magnetization is generated, exceeding the in-plane EE. For the latter term, the transverse magnetization is the only response allowed, and it is enhanced when the chemical potential is close to the crossing of the electronic bands due to RSOC. In both cases, it is remarkable that out-of-plane magnetization is obtained in response only to in-plane fields. All these features can be exploited to fully characterize the altermagnetic nature of the electron gas at the interface. Indeed, the behaviour of the magnetization with the filling can be directly related to the presence of the spin-orbit, DM interaction and band crossings. On the other hand, the uncovered anomalous non-linear altermagnetic Edelstein responses may be used experimentally to induce efficient spin-to-charge conversion at altermagnetic interfaces leading to novel electronic and spintronic applications.
@@
Acknowledgments — We thank Oleg Janson and Carmine Autieri for fruitful discussions. We thank Ulrike Nitzsche for technical asistance. M.T. acknowledges financial support from “Fondazione Angelo Della Riccia”. This work was financed by Horizon Europe EIC Pathfinder under the grant IQARO number 101115190. R.C. and F.R. acknowledge funding from Ministero dell’Istruzione, dell’Università e della Ricerca (MIUR) for the PRIN project QUESTIONS (Grant No. PRIN P2022KWFBH) and STIMO (Grant No. PRIN 2022TWZ9NR). This work received support from the PNRR MUR project PE0000023-NQSTI (TOPQIN and SPUNTO). This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Sonderforschungsbereich SFB 1143, Grant No. YE 232/2-1, and under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, Project IDs No. 390858490 and No. 392019).
Contributions — R.C., M.T., I.G, and J.v.d.B. conceived the ideas and designed the research; M.T., I.G. and C.G. performed the transport calculations; J.I.F. performed DFT calculations; all the authors interpreted and analized the results; M.T., C.G. R.C., J.I.F. and J.v.d.B. wrote the first version of the manuscript with input from all authors; all the authors reviewed the manuscript; R.C. and J.v.d.B. supervised the study.
Appendix A Scaling in the electric and magnetic field
The effect of RSOC and the magnetic field to the altermagnetic system, can be undestood on the basis of a heuristic argument based on the accumulated spin considering only four points of the Fermi surface. The Fermi surface generated from the pure altermagnetic Hamiltonian is
$H=\frac{k^{2}}{2m}+\gamma(k_{x}^{2}-k_{y}^{2})\sigma_{z},$ (6)
which has the form of two spin-polarized orthogonal ellipses centered around $\Gamma$. Let us consider the four points at the end of the major axis of the ellipses in the momentum space ±$\pm K_{x}=(\pm K,0)$ and ±$\pm K_{y}=(0,\pm K)$, where $K$ is a large enough quasi-momentum, such as $\gamma K\gg\alpha$ and $\gamma K^{2}\gg h_{y}$, where $\alpha$ and $h_{y}$ are the RSOC parameter and the magnetic Zeeman field that will act as a perturbation. Let us consider the perturbation due to the RSOC and the magnetic field
$V=\alpha(k_{y}\sigma_{x}-k_{x}\sigma_{y})+h_{y}\sigma_{y},$ (7)
the spin along $z$ is therefore
$s_{z}=\frac{\gamma(k_{x}^{2}-k_{y}^{2})}{\sqrt{\gamma^{2}(k_{x}^{2}-k_{y}^{2})^{2}+\alpha^{2}(-k_{x}+h_{y})^{2}+\alpha^{2}k_{y}^{2}}}.$ (8)
We consider the electric field applied to the system by taking into account the shift in the momentum $\delta k$ due to the equation of motion, i.e., $E_{x}=\frac{\tau\delta k}{e}$, where $\tau$ is the characteristic scattering time of the system.
Therefore, the total spin along $z$ is {dmath} S_z=s_z(K+δk,0)+ s_z(-K+δk,0)+ s_z(δk,K)+ s_z(δk,-K), which we expand to first order in $h_{y}$ and to second order in $\delta k$, so that
$S_{z}=\frac{3\alpha^{4}\gamma K^{4}}{\left(K^{2}\left(\alpha^{2}+\gamma^{2}K^{2}\right)\right)^{5/2}}\delta k^{2}-\frac{2\left(\alpha^{2}\gamma\left(\alpha^{2}+4\gamma^{2}K^{2}\right)\right)}{K(\alpha^{2}+\gamma^{2}K^{2})^{5/2}}h_{y}\delta k=F\,E_{x}^{2}+G\,h_{y}E_{x},$ (9)
where $F$ and $G$ are two constants related to microscopic parameters. From this simple consideration one expects that in the absence of magnetic field the linear anomalous out-of-plane EE vanishes, while it persists as a second-order response in the electric field.
Appendix B Parameter fit for RuO2 bilayer
@@
Figure 6: Fit of the bands crossing the Fermi level nearby $\Gamma$ along the direction $\Gamma Y$.
Here we extract the parameter $\gamma$ of Hamiltonian (1) from DFT data of RuO2 bilayer. In Fig. 6 we show the DFT data of RuO2 and the fit of the band structure for the two bands crossed by the Fermi level. We fitted these two bands using the quadratic expression $E_{i}=a_{i}k_{x}^{2}+b_{i}k_{x}+c_{i}$, where we truncated the DFT bands in order to minimize vertex location of the two parabolas, therefore by minimizing $\bar{k}_{i}=|\frac{b_{i}}{2a_{i}}|$. We obtain from the fitting procedure
@@
By comparison with the analytical expression of the eigenvalues of Eq. (1) with $\alpha=0$, $\mathbf{h}=0$, and $k_{y}=0$, which are
$E_{\pm}=\left(\frac{1}{2m}\pm\gamma\right)k_{x}^{2},$ (11)
we obtain the value of $\gamma$ by the ratio
$\frac{(a_{+}-a_{-})}{(a_{+}+a_{-})}=2m\gamma=0.571.$ (12)
For $m=1$ we obtain $\gamma\approx 0.3$. In absolute value $m=0.152$ $\text{eV}^{-1}\text{\AA}^{-2}$ (which is $1.2m_{e}$, where $m_{e}$ is the electron mass) and $\gamma=1.881$ $\text{eV}\text{\AA}^{2}$. The sign of $\gamma$ cannot be determined by band comparison alone, but can be fixed by comparison with the spin patterns throughout the Brillouin zone. From Fig. 7, we see that the lower band along the $\text{Y}_{0}\Gamma$ path is polarized with spin down, indicating that $\gamma>0$.
Figure 7: Spin projection along $z$ direction of RuO2 bilayer.
Appendix C Behaviour of the response with parameters
This section details the dependence of the Edelstein responses on the RSOC parameter ($\alpha$) and the Zeeman coupling ($h_{y}$). In Fig. 8 we show the in-plane and out-of-plane susceptibility for two values of $\alpha$. Within this energy window, the response is clearly linear in $\alpha$. We can understand this result invoking a simple energy scale analysis; the RSOC term is larger than the altermagnetic one for $k\lesssim\alpha/\gamma\sim 10^{-2}$-1, and therefore for energies $\varepsilon\lesssim\alpha^{2}/2m\gamma^{2}\sim 10^{-3}$ eV. On the other hand, below these energies, the dominant energy scale is the weak ferromagnetic coupling $h_{y}$. Therefore, across the entire energy range, the RSOC remains a perturbation. The response vanishes for small RSOC.
Figure 8: (a) $\chi_{xy}$ and and (b) $\chi_{xz}$ as a function of the chemical potential for $\alpha=13\text{ meV \AA}$ and $4\alpha=52\text{ meV \AA}$, together with the value of (a) $4\chi_{xy}$ and (b) $4\chi_{xz}$, using $\gamma=1.881$ eV Å2, $h_{y}=40$ meV.
@@ Appendix C Behaviour of the response with parameters
In Fig. 9, we show the behaviour of the EE for different $h_{y}$. The key insight is that the in-plane response is even in $h_{y}$, remaining identical for $h_{y}\to-h_{y}$, while the out-of-plane response is odd, changing its sign. Both results are easily understood; the inversion of the magnetic field $h_{y}\to-h_{y}$ can be canceled by a corresponding change in sign $k_{x}\to-k_{x}$, which leaves the eigenvalues of the Hamiltonian unchanged. With such a transformation, the group velocity $v_{x}\to-v_{x}$, and the same goes for the in-plane spin $s_{y}\to-s_{y}$, while the out-of-plane spin, determined by the altermagnetic term, which is even in $k_{x}$, remains unchanged $s_{z}\to s_{z}$. Therefore, from the definition of the EE susceptibility, we see that the in-plane response is even, while the out-of-plane response is odd in $h_{y}$. Correspondingly, the out-of-plane response, which vanishes for $h_{y}\to 0$, must have a linear behavior $\chi_{xz}\propto h_{y}$ for very small fields. In Fig. 10 we show $\chi$ as a function of the Zeeman coupling for a fixed value of the chemical potential $\mu/h_{y}\gg 1$, in order to show the scaling in the saturation region.
Figure 9: (a) $\chi_{xy}$ and (b) $\chi_{xz}$ as a function of the chemical potential for $\alpha=52$ meV Å and different values of the in-plane Zeeman field $h_{y}$.
Second-order Edelstein susceptibility— The ratio between of the RSOC and the altermagnetic energy scales is determined by $|\alpha|/|k_{F}\gamma|$, with $k_{F}$ to be the Fermi momentum. It is a measure for their competition, which is strongest when $|\alpha|/|k_{F}\gamma|\simeq 1$. For the value $|\alpha|/|k_{F}\gamma|\simeq 1$ we observe that the second-order anomalous susceptibility has a peak. In Fig. 11 we show the second-order anomalous Edelstein susceptibility for different values of $\gamma$. It tends to peak at a small value of the chemical potential close to the band crossing at $\Gamma$ (see Figure 2). The peak is substantially broadened when $|\gamma k_{F}|>|\alpha|$. Indeed the peak-shape of the second-order response is lost for small values of $\gamma$ ($\gamma=0.2$ eV Å2). Within the considered range of $\mu$ values, the maximum explored Fermi momentum corresponds to $k_{F}=0.22$-1, so that $|\alpha|>|\gamma k_{F}|$. However, since the non-relativistic spin splitting is expected to be larger (as seen from DFT calculation), the peak-shape feature might be used as a probe of the band crossing due to the Rashba coupling. To see why the non-monotonicity arises, it is useful to look at the integrand in Eq. (20) by moving the chemical potential. We show in Fig. 12 the integrand for different chemical potentials, using as benchmark values $\gamma=1.881$ eV Å2 and $\alpha=52$ meV Å . We can see that, as long as the chemical potential is close to the band crossing, the inner band contribution is stronger than the outer one. Moreover, the integrand structure is peaked were the avoided crossings are located. For higher filling the two contributions compensate and the total integral approaches to zero.
@@
Figure 10: Edelstein susceptibility behaviour with Zeeman coupling for $\alpha=52$ meV Å and for fixed chemical potential $\mu=1.05$ eV and $\gamma=1.881$ eV Å2. Lines are guides to the eye.
@@
Figure 11: Second order Edelstein susceptibility as a function of the chemical potential for $\alpha=52$ meV Å and different values of $\gamma$. The dashed vertical line marks the value of $\mu$ used for Fig. 13(a).
In order to quantify the effect of the maximum efficiency of conversion, in Fig. 13(a) we show how the response is behaving by increasing $\gamma$ at a fixed chemical potential far away from the peaked region, i.e., $\mu=0.011$ that is the vertical dashed line in Fig. 11, and in Fig. 13(b) the maximum value of the peak for each $\gamma$; in both cases we observe a non-monotonic behaviour.
Figure 12: (a-d) Contour plots of the integrand $U=\sum_{\nu}\nu\partial_{k_{x}}s_{z}v_{x}^{\nu}\delta(\varepsilon^{\nu}-\mu)$ in Eq. (5) for different values of $\mu$ corresponding to the dashed vertical lines of the bottom panel. The benchmark parameters have been chosen as $\gamma=1.881$eV Å2 and $\alpha=52$ meV Å. The Dirac delta function has been substituted with the derivative of the thermal Fermi function, with a temperature of $T=33$ K, corresponding to $T=2.8$ meV.
Figure 13: (a) Dependence of the second-order Edelstein susceptibility on the altermagnetic coupling $\gamma$ (expressed in eV Å2), for $\alpha=52$ meV Å and for fixed chemical potential $\mu=11$ meV, which corresponds to the dashed vertical line in Fig. 11. (b) Maximum of second-order Edelstein susceptibility versus the Zeeman coupling for $\alpha=52$ meV Å. Lines are guides to the eye.
@@
Finally, to quantify the effect of the RSOC, we show in Fig. 14 the dependence on $\alpha$ of the second-order Edelstein susceptibility. As expected from Eq. (9), for a fixed chemical potential the susceptibility is an even function of $\alpha$, showing also a sign change for $\alpha$ close to zero.

### v3

@@ III Calculation of non-linear Edelstein responses
Having established that, due to the interplay of RSOC with AM symmetry allows anomalous non-linear EEs, we now determine them in an experimentally relevant parameter range. For the effective in-plane field $h_{y}$ resulting from weak ferromagnetism, we choose $h_{y}=40$ meV as benchmark value [38]. The first effect of the in-plane Zeeman coupling is to lift the degeneracy at $\Gamma$, as shown in Fig. 3(a), in contrast with the simple Rashba system, for which the in-plane magnetic field can only move the degeneracy point. Moreover, as mentioned above, it breaks the symmetry of the system both in the shape of the Fermi surface and in the chiral structure of the spin, as one can see in Fig. 3(b). The resulting Edelstein susceptibility tensor $\chi$ is shown in Fig. 3(c). While $\chi_{xx}$ vanishes, $\chi_{xy}$ depends non-monotonically on chemical potential and reaches its maximum after occupying the first band. Compared to the case $h_{y}=0$, $\chi_{xy}$ grows slower to its maximum, but reaches a value similar to the situation with pure RSOC. Moreover, the in-plane magnetic field is able to induce a substantial out-of-plane anomalous transversal response $\chi_{xz}$ due to the large altermagnetic coupling, with the transversal component reaching larger values than the in-plane response, even larger than the canonical linear RSOC situation.
@@
The non-monotonic behaviour of $\chi_{xz}$ as a function of $\mu$ derives from the competition and complex interplay of the different interactions. At low filling, even an infinitesimal $h_{y}$ breaks the $\mathcal{T}C_{4z}$ symmetry, leading to a characteristic threshold behavior in the response. Increasing $\mu$, the out-of-plane response is first slightly suppressed and later saturates to a constant value in the altermagnetic-dominated regime. Crucially, all three ingredients are required to exhibit the out-of-plane response: the altermagnetic interaction enforces an out-of-plane magnetization, the RSOC induces the spin-momentum locking, and the weak ferromagnetism breaks its otherwise symmetric pattern required for this non-linear Edelstein response.
@@ IV Summary and outlook
Using a continuum description for AMs based on ab-initio electronic band structure calculations, we have shown that the interplay of altermagnetic coupling with an interfacial RSOC gives rise to the appearance of out-of-plane spin polarizations throughout the Brillouin zone via non-linear EEs. While the in-plane linear response can be strongly suppressed with the band filling, the non-linear Edelstein susceptibilities exhibit strong non-monotonic behaviour as a function of the chemical potential. The non-linear responses we determined involve a term that depends on the in-plane electric and the magnetic fields, and a second-order term in the electric field. For the former term, an anomalous out-of-plane magnetization is generated, exceeding the in-plane EE. For the latter term, the out-of-plane magnetization is the only response allowed, and it is enhanced when the chemical potential is close to the crossing of the electronic bands due to RSOC. In both cases, it is remarkable that out-of-plane magnetization is obtained in response only to in-plane fields. All these features can be exploited to fully characterize the altermagnetic nature of the electron gas at the interface. Indeed, the behaviour of the magnetization with the filling can be directly related to the presence of the spin-orbit, DM interaction and band crossings. To isolate the essential physics, we focused on a minimal model capturing the local interplay between altermagnetic exchange and Rashba spin-orbit coupling near the $\Gamma$ point. This approach reveals the core mechanism underlying the anomalous Edelstein response in altermagnets. Other band crossings at high-symmetry points can be described within the same framework (see Appendix E), and while a full multiband analysis may introduce quantitative differences, it does not alter the fundamental physical picture. The uncovered anomalous non-linear altermagnetic Edelstein responses may be used experimentally to induce efficient spin-to-charge conversion at altermagnetic interfaces leading to novel electronic and spintronic applications.
V Acknowledgments
@@
We thank Oleg Janson and Carmine Autieri for fruitful discussions. We thank Ulrike Nitzsche for technical assistance. M.T. acknowledges financial support from “Fondazione Angelo Della Riccia”. This work was financed by Horizon Europe EIC Pathfinder under the grant IQARO number 101115190. R.C. and F.R. acknowledge funding from Ministero dell’Istruzione, dell’Università e della Ricerca (MIUR) for the PRIN project QUESTIONS (Grant No. PRIN P2022KWFBH) and STIMO (Grant No. PRIN 2022TWZ9NR). This work received support from the PNRR MUR project PE0000023-NQSTI (TOPQIN and SPUNTO). This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Sonderforschungsbereich SFB 1143, Grant No. YE 232/2-1, and under Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter – ct.qmat (EXC 2147, Project IDs No. 390858490 and No. 392019).
Appendix A Spin-texture analysis
In this section we analyze in detail the spin-texture of the Hamiltonian including also the Zeeman coupling
$\hat{H}=\frac{k^{2}}{2m}+\gamma(k_{x}^{2}-k_{y}^{2})\hat{\sigma}_{z}+\alpha\hat{z}\cdot(\vec{\sigma}\times\mathbf{k})+h_{y}\hat{\sigma}_{y}-\mu.$ (6)
The eigenvalues of this $2\times 2$ matrix are easily obtained, as well as the direction of the spin quantization axis
$\varepsilon^{\pm}=\frac{k^{2}}{2m}\pm\sqrt{(h_{y}-\alpha k_{x})^{2}+\alpha^{2}k_{y}^{2}+\gamma^{2}(k_{x}^{2}-k_{y}^{2})^{2}},$ (7)
$\mathbf{s}=\frac{(\alpha k_{y},\,h_{y}-\alpha k_{x},\,\gamma(k_{x}^{2}-k_{y}^{2}))}{\sqrt{(h_{y}-\alpha k_{x})^{2}+\alpha^{2}k_{y}^{2}+\gamma^{2}(k_{x}^{2}-k_{y}^{2})^{2}}}.$ (8)
For $\gamma=0$, the band structure exhibits a crossing at $\mathbf{k}=(h_{y}/\alpha,\,0)$. However, the spin texture is tilted along the $y$-direction, distorting the in-plane chirality. When $\gamma\neq 0$, the $s_{z}$ component of the spin becomes an odd function of $\mathbf{k}$, as the parity symmetry is broken by the denominator of $\mathbf{s}$.
Notably, for $\alpha=0$, the presence of $h_{y}$ alone does not break the even symmetry of $s_{z}$ (i.e. $s_{z}(\mathbf{k})=s_{z}(-\mathbf{k})$). This highlights the need for all three ingredients—Rashba spin-orbit coupling ($\alpha$), Zeeman field ($h_{y}$), and anisotropic coupling ($\gamma$)—to generate an out-of-plane Edelstein effect. In Fig. 5 we show the values of $s_{y}$ and $s_{z}$ as a function of $k_{x}$ for $k_{y}=0$ for $\alpha\neq 0$ and by varying $\gamma$ and $h_{y}$.
Figure 5: $s_{z}$ (a) and $s_{y}$ (b) as a function of $k_{x}$ and $k_{y}=0$ with benchmark values of $\alpha\neq 0$ and by varying $\gamma$ and $h_{y}$. The values of $\alpha$, $\gamma$ and $h_{y}$ have been chosen to make the plot clearer.
In case of a pure Rashba coupling only $s_{y}$ is present and is constant with a change of sign at $k_{x}=0$, while when a Zeeman field is present along the line $k_{y}=0$ the spin is constan. When the altermagnetic coupling is turned on, the even (odd) symmetry in $k_{x}$ is present for $s_{z}$ ($s_{y}$), which is broken while $h_{y}\neq 0$.
Appendix B Ab-initio RuO2 bilayer band structure and parameter fit
Here we estimate the strength of the altermagnetic coupling $\gamma$ by explicit comparison with the local band structure of a RuO2 bilayer. Based on the bulk crystal structure, we construct a bilayer oriented along the $c$-axis of RuO2 bulk, containing two Ru ions related by fourfold rotoinversion symmetry. The altermagnetic bilayer can be represented with the space group Cmm2 (No. 35), with a unit cell having $a=6.35$ Å. For this structure, we perform density-functional calculations in the generalized gradient approximation based on the FPLO code [29]. We use a $k$-points mesh having $16\times 16\times 1$ subdivisions along with a tetrahedron method for Brillouin zone integrations. Figure 6 shows the band structure in the absence of RSOC. We apply our effective two-band model to describe the two bands closest to the Fermi energy at $\Gamma$.
Refer to captiond)
@@ Appendix B Ab-initio RuO2 bilayer band structure and parameter fit
Figure 6: (a-c) Bilayer of RuO2 and its electronic structure density-functional calculations. (a) Crystal structure. (b) Band structure along high symmetry points indicated in (c). (d) Fit of the bands crossing the Fermi level nearby $\Gamma$ along the direction $\Gamma Y$.
The parameter $\gamma$ and $m$ of Hamiltonian (6) (for $\alpha=0$ and $h_{y}=0$) can be extracted from DFT data of RuO2 bilayer. We note here that our model has a spin-ordering rotated of $45^{\circ}$ with respect to the RuO2, which simplifies the analysis of the Edelstein susceptibility tensor. In Fig. 6 we show the DFT data of RuO2 and the fit of the band structure for the two bands crossed by the Fermi level. The two bands are well fit by
@@
By comparison with the analytical expression of the eigenvalues of Hamiltonian (6) we obtain $m=0.152$ eV-1 Å-2 and $\gamma=1.881$ $\text{eV}\text{\AA }^{2}$.
Appendix C Parametric dependence of the Edelstein response
@@
This section details the dependence of the Edelstein responses on the RSOC parameter ($\alpha$) and the Zeeman coupling ($h_{y}$). In Fig. 7, we show the behaviour of the EE for different $h_{y}$. The in-plane response is even in $h_{y}$, while the out-of-plane response is odd in $h_{y}$. The inversion of the magnetic field $h_{y}\to-h_{y}$ can be canceled by a corresponding change in sign $k_{x}\to-k_{x}$, which leaves the eigenvalues of the Hamiltonian unchanged. With such a transformation, the group velocity $v_{x}\to-v_{x}$, and the same goes for the in-plane spin $s_{y}\to-s_{y}$, while the out-of-plane spin, determined by the altermagnetic term, which is even in $k_{x}$, remains unchanged $s_{z}\to s_{z}$. Therefore, from the definition of the EE susceptibility, we see that the in-plane response is even, while the out-of-plane response is odd in $h_{y}$. Correspondingly, the out-of-plane response, which vanishes for $h_{y}\to 0$, must have a linear behavior $\chi_{xz}\propto h_{y}$ for very small fields. In Fig. 7 we show $\chi$ as a function of the Zeeman coupling for a fixed value of the chemical potential $\mu/h_{y}\gg 1$, in order to show the behvaviour of the $\chi$ in the saturation region.
Refer to caption(c)
@@
Figure 7: (a) $\chi_{xy}$ and (b) $\chi_{xz}$ as a function of the chemical potential for $\alpha=52$ meV Å and different values of the in-plane Zeeman field $h_{y}$. (c) Edelstein susceptibility behaviour with Zeeman coupling for $\alpha=52$ meV Å and for fixed chemical potential $\mu=1.05$ eV and $\gamma=1.881$ eV Å2. Lines are guides to the eye.
Second-order Edelstein susceptibility— The ratio between of the RSOC and the altermagnetic energy scales is determined by $|\alpha|/|k_{F}\gamma|$, with $k_{F}$ to be the Fermi momentum. It is a measure for their competition, which is strongest when $|\alpha|/|k_{F}\gamma|\simeq 1$. For the value $|\alpha|/|k_{F}\gamma|\simeq 1$ we observe that the second-order anomalous susceptibility has a peak. In Fig. 8 we show the second-order anomalous Edelstein susceptibility for different values of $\gamma$. It tends to peak at a small value of the chemical potential close to the band crossing at $\Gamma$ (see Fig. 2 of main text). The peak is substantially broadened when $|\gamma k_{F}|>|\alpha|$. Indeed the peak-shape of the second-order response is lost for small values of $\gamma$ ($\gamma=0.2$ eV Å2). Within the considered range of $\mu$ values, the maximum explored Fermi momentum corresponds to $k_{F}=0.22$ Å-1, so that $|\alpha|>|\gamma k_{F}|$. However, since the non-relativistic spin splitting is expected to be larger (as seen from DFT calculation), the peak-shape feature might be used as a probe of the band crossing due to the Rashba coupling.
Refer to caption(a)
Refer to caption(b)
@@
Figure 8: (a) Second order Edelstein susceptibility as a function of the chemical potential for $\alpha=52$ meV Å and different values of $\gamma$. (b) Second-order Edelstein susceptibility $\tilde{\chi}_{xxz}$ as a function of $\alpha$ for $\mu=0$ and $\gamma=1.881$ eV Å2. The line is a guide to the eye.
@@
Finally, to quantify the effect of the RSOC, we show in Fig. 8 the dependence on $\alpha$ of the second-order Edelstein susceptibility. For a fixed chemical potential the susceptibility is an even function of $\alpha$, showing also a sign change for $\alpha$ close to zero.